Compute the cost of not taking the following cash discounts. (Use a 360-day year. Do not round intermediate calculations. Input your final answers as a percent rounded to 2 decimal places.) Cost of Lost Discount a. 214, net 45 b. 219, nct 60 c. 4/19, net 45 d. 3/13, net 130
Answer #1

Calculation of cost of not taking cash discount

a)2/14, net 45 it means the bussiness wil pay full amount of invoice i.e 2% more than its needs to, if it wait for the full 30 days period to elapse before making payments instead of paying within 10 day period,so there is cost to bussiness of not taking the discount.

The formula is= The percentage of discount/(100-percentage of discount)×( 360/number of days between early payment date and normal due date)×100

The percentage of discount=2

number of days between early payment date and normal due date= 45-14=31

cost of not taking discount=2/(100-2)×360/31

=(2/98)×(360/31)×100=23.69%

(B)2/19,net 60

The formula is= The percentage of discount/(100-percentage of discount)× (360/number of days between early payment date and normal due date)×100

The percentage of discount=2

number of days between early payment date and normal due date= 60-19=41

cost of not taking discount=2/(100-2)×360/41

=(2/98)×(360/41)×100=17.91

C)4/19 ,net 45

The formula is= The percentage of discount/(100-percentage of discount)× 360/number of days between early payment date and normal due date×100

The percentage of discount=4

number of days between early payment date and normal due date= 45-19=26

cost of not taking discount=4/(100-4)×360/26

=(4/96)×(360/26)×100=57.69%

D)3/13 net 130

The formula is= The percentage of discount/(100-percentage of discount)× 360/number of days between early payment date and normal due date×100

The percentage of discount=3

number of days between early payment date and normal due date= 130-13=117

cost of not taking discount=3/(100-3)×360/117

=(3/97)×(360/117)×100=9.56%